Integrate the following function with respect to t

 

f(t) = 0,2e^−0.2t


a.− e^−0.2t + c


b. e^−0.2t + c


c.−o.4e^−0.2t


d.−o.4e^−0.2t + c



1
Expert's answer
2022-05-03T09:18:47-0400

"\int0.2\,e^{-0.2\,t}dt=-\int\,e^{-0.2\,t}d(-0.2t)=-e^{-0.2\,t}+C," "C\in{\mathbb{R}}". We used a formula for a well-known integral "\int e^x dx" and a substitution: "z=-0.2\,t".

Answer: a. "-e^{-0.2\,t}+C".